OpenStudy (anonymous):

a costume designer made 10 costumes that each used 5.8 yards of fabric if she had 238.2 yards of fabric to start with about how much fabric does the costume designer have left? write an equation estimate the amount of fabric left explain answer

OpenStudy (tapha):

ok

OpenStudy (tapha):

180.2

OpenStudy (tapha):

got it

OpenStudy (anonymous):

amount of fabric left=238.2-10x5.8=238.2-58=180.2

OpenStudy (anonymous):

thankyou